Group Selection Page
HELP
Controls Group
Data Analysis Group
ITDM
WP76
DET
Vacuum
EEE
WP78
WP74
FXE
SCS
SPB
MID
HED
SQS
SXP
Sample Environment
Operation
PSPO
XO
TS
Migrated2Zulip
|
Technical integration and coordination
SASE1 Status and SASE3 Readiness
|
Joint Operation and Readiness Logbook, Contact: T.Haas, A. Violante, G. Wellenreuther |
Not logged in |
 |
|
Message ID: 539
Entry time: 27 Jun 2018, 14:55
|
Author: |
Adriano Violante |
Location: |
General |
Severity: |
Information |
Subject: |
Status report after dispatch meeting 13:00 |
|
|
SPB
- Electronic and instrumentation will remain OFF until the electricity comes back tomorrow (will be announced by TS on the logbook)
- MIrrors installation will start next week
- Mirrors movement test to happen, then Loop 4 and 1 to be updated
- In order to make the calibration of the motors easier after the powering up, they should not be moved before.
SCS
- New cables arrived. Testing ongoing.
- Loop 1,2,3, up. 1 has to be updated. Coordination AE/MTeichmann ongoing
- Loop4 and 5 being prepared by AE
AE
- XTD6 DPS and MCP being commissioned.
- SCS issues with Beckhoff modules being fixed.
- AE and CAS to agree on interlock reset procedure.
CAS
- Karabo 2.2.4. will be deployed in the SASE3 tunnels tomorrow at 10:00. It will take approx. 2 hours.
|